# LakeVilla: Multi-Table Transactions for Lakehouses

### v1

@@ 4.2. Isolation
Base OTFs: Serializability (Single-Query, Single-Table). OTF formats guarantee serializability limited to single-query and single-table transactions. We observe that all OTF transactions define a sequential order among their operations. Further, OTFs like Delta Lake define a serial order between all log files in their metadata layer (Armbrust et al., 2020). Since OTFs only feature single-query and single-table transactions, all transaction operations are contained within a single log file, providing serializability or snapshot isolation (Databricks, 2024; Jain et al., 2023). As a result, the execution order of all transactions equals a serial execution. Their operations can be ordered transitively to form a serial history between all operations on the same table.
LV [R]: Serializability for single tables. LV [R] extends the guarantee of the base OTFs by adding atomic markers placed during a transaction’s start. Hence, all distinct transactions operating on the same table initially get a unique version. Using atomic object store operations guarantees uniqueness among the assigned versions for each transaction. This defines the version a transaction will use during its execution. Further, LV [R]’s marker technique guarantees that all transaction operations are contained in the same log file of the reserved version. This effectively defines a sequential ordering of transactions and their operations on the same table. Like in the base OTFs, transactions are sequentially ordered by the OTF structure and operations within their transaction. Hence, because the markers ensure correct positioning of all operations within the metadata layer, they uphold the structure of the base OTFs. At the same time, markers make it possible to put any number of operations into the correct position, dropping the original OTF’s limitation on single-query transactions.
LV [R, CT]: Serializability. With LV [R]’s shift to a table-level focus, LV [CT] guarantees repeatable reads of the entire LH. Like LV [R], LV [CT] places atomic markers in a table to claim a version for a transaction. However, it extends this model to claim individual table versions in all accessed tables. The defined deadlock detection model and marker shifts guarantee that the same order among related transactions is upheld across tables. A client accessing a new table will detect any previous markers within. This defines the order in which the transaction depends on all other transactions of the discovered markers. If these dependencies form a cycle or timeout, LV [CT]’s marker shift inverts the previous dependency in all other tables, ensuring it is upheld. If the newly found dependencies do not violate previously found ones, no marker shift is required, and local dependencies become globally enforced on all tables. Hence, because a client must redo some of its operations after a marker shift, LV [R, CT] guarantees serializability globally.
LV [CT]: Repeatable Reads. However, without LV [R], LV [CT] enables phantom reads after a marker shift. A transaction executing a query before a marker shift may receive a different result set for ranges afterward. Caused by the marker shift, the table versions between transactions were reversed, and newer changes that might have affected the result set were adapted. However, by remembering previous reads before the marker shift, reusing the old read versions, and checking for compatibility during commit, LV [CT] can prevent non-repeatable reads. Hence, LV [CT] guarantees the repeatable read isolation level.
LV [I]: Serializability (LV [I] only). LV [I] provides the isolation level of serializability. Its global version log defines a total order between all committed transactions. Each transaction must enter its atomic validation phase and check its changes using R1 – R3. It can only succeed if no other committed transaction changed its write set; otherwise, it must abort. Hence, the order of successfully validated transactions gives a sequential global order. Due to the atomic design of the validation phase, LV [I] guarantees that previously validated operations will be ordered before and after the current transaction. Similarly, as the global version log is a global layer for the entire LH, it defines relations between all transactions, giving Serializability. Hence, LV [I]’s ordering is not limited to table level. If transactions only check R2 – R3 during validation, we get Snapshot Isolation as we only check write sets. However, concurrent operations not using LV [I] ignore the global layer. To avoid inconsistencies, LV [I] uses the hidden version to adapt to such changes and still perform its validation, guaranteeing Serializability/Snapshot Isolation among LV [I] transactions.
4.3. Experimental Validation
@@
Table 3. Shortened Hermitage Results (multi-query) for Spark using Delta Lake (Spark [DL]) or Iceberg (Spark [IB]) and LakeVilla [LV]. Anomalies described with ”-MT” are multi-table versions of the original scenario. We tested all anomalies and show the ones with the biggest differences here.
@@
G0-MT (Write Cycles) $faThumbsOUp * \faThumbsODown \faThumbsOUp \faThumbsOUp$
G1a (Aborted Reads) $faThumbsODown \faThumbsOUp \faThumbsOUp \faThumbsOUp$
G1b (Interm. Reads) $faThumbsODown \faThumbsOUp \faThumbsOUp \faThumbsOUp$
G1c-MT (Circ. information flow) $faThumbsODown * \faThumbsODown \faThumbsOUp \faThumbsOUp$
OTV (Obs. Txn. Vanishes) $faThumbsODown \faThumbsOUp \faThumbsOUp \faThumbsOUp$
P4 (Lost Update) $faThumbsODown \faThumbsOUp \faThumbsOUp \faThumbsOUp$
Concurrent Reads (Linearizability) - - $faThumbsODown \faThumbsOUp$
[…]
@@
To provide further validation of LV’s guarantees, we experimentally check Spark and our prototype using LV [R], LV [R, CT], and LV [R, CT, I] for the anomalies described by Hermitage (Kleppmann, 2014). A selection of the results is shown in Tab. 3 with additional multi-table versions of the same anomaly (marked with -MT). The original Delta Lake (Spark [DL]) and Iceberg (Spark [IB]) were included despite their restriction on single-query transactions, which led them to fail most multi-query anomalies. An exception is G0-MT (write cycles), avoided by their limitations and thus marked uniquely (G0-MT). In contrast, LV [R] prevents table-level anomalies like lost updates (P4), aborted reads (G1a), or Intermediate Reads (G1b) and adapts to aborted/crashed transactions (OTV). When adding LV [CT], the design can resolve conflicts across tables like Write Cycles (G0-MT) and Circular information flow (G1c-MT). Further, LV [I]’s global isolation guarantees concurrent transactions to observe globally consistent versions across tables. In summary, these results confirm our targeted guarantees.
5. Evaluation
@@
Overview. We start by analyzing the different phases of the LV Features and comparing them to Spark’s Delta Lake (Foundation, 2018; Projects, 2024a). We investigate their effects under concurrent workloads using YCSB [^1] YCSB-LH available: https://github.com/goetztj/YCSB-LH (Cooper et al., 2010; Ren et al., 2024) and test different feature combinations based on the cloud analytics benchmark (CAB) (van Renen and Leis, 2023). We conclude by running LV on selected analytical queries from TPC-DS and estimating the added overhead for systems like Spark.
5.1. Experimental Setup
@@
Object Store: MinIO. For all experiments, we use MinIO (Foundation, 2024b) on a cloud instance with an Intel(R) Xeon(R) Platinum 8260 CPU with 48 cores at 2.40GHz with 524GB of RAM and 5TB of SSD storage running Linux. MinIO provides strict consistency (Sarrel, 2024), conditional writes (Amazon Web Services, 2014), and analytical features for the experiments. It is accessible through the AWS S3 interface and behaves like AWS’s offering.
@@
Spark: Delta Lake. Spark 3.5.3 runs with Hadoop 2.12, Delta Lake 3.0.0 (checkpoint interval 10), and Hadoop AWS 3.3.4. We minimized caching to observe the OTFs’ effects on the system. Our HIVE 3.0.0 catalog and its PostgreSQL backend run in two separate docker containers. Spark uses Java 11 and runs distributed on 9 machines with the same setup as above. The coordinator runs on one server with specifications like our MinIO server and shares its resources with the catalog store and benchmark drivers to reduce the networking effects between these components. For workers, we configured up to two instances on eight servers using an Intel(R) Xeon(R) Platinum 8260 CPU at 2.40GHz with 8 cores each. Each worker was assigned 64GB RAM and eight cores.
5.2. Phases of OTFs
@@
Figure 5. Phases for read and write workloads of a single transaction for YCSB queries.
[…]
[Phases for read and write workloads of a single transaction for YCSB queries.]Phases for read and write workloads of a single transaction for YCSB queries. The prototype’s reads are comparable to Spark, while Spark’s writes spend more time on the OTF version assignment.
Comparison of Spark and LakeVilla. We implemented our LV prototype as a separate C++ client. Thus, we first compare the differences between it and other systems. For this purpose, we ran YCSB sequentially with Spark using Delta Lake (Spark [DL]) and LV [DL]. If not stated otherwise, we use the default YCSB of 1000 operations on 1000 entries with row sizes of 1KB (Cooper et al., 2010). We logged the different phases of both implementations and calculated both the average latency for reads and writes, and how much each phase contributes to the overall latency. We expect both systems to display similar behavior. We show the results in Fig. 5. Spark [DL] and LV [DL] indeed behave similarly for reads. They spend most of their time accessing and reading the metadata layer of the YCSB table (95% and 93%, respectively) and the rest to load the requested data.
@@
The different LV Features. Next, we evaluate the impact of each LV feature on latency. We use the same setup as before and show our results in Fig. 5(a). First, we analyze the overhead of placing LV [R] ’s and LV [CT] ’s markers (technique #1). We hypothesize that it affects the latency, especially for reads. In Fig. 5(a), writes show minimal impact as they align with metadata layer operations in the original OTF (LV [DL]). Further, freeing those markers is fused with the write transaction’s log placement, compensating its effect as seen in Fig. 5(a). In contrast, reads experience a 20% latency increase due to added OTF synchronization, with an additional impact when we free the markers post-read. This indicates that LV [R] and LV [CT] are better suited for write operations.
@@
Next, we analyze the impact of LV [R] ’s sublog (technique #2) and LV [CT] ’s deadlock detection (technique #3) on latency using the same setup. Because both techniques are expected to work concurrently alongside the standard operations, we expect minimal impact on latency, which is also confirmed by Fig. 5(a). The time spent by the prototype on the actual operation in LV [R] and LV [CT] matches the baseline (LV [DL]). LV [R] ’s sublog generation is distinct and does not impact the baseline operation. Further, LV [CT] ’s deadlock detection does not detect problems as YCSB uses only one table. Hence, under non-concurrent workloads, neither technique impacts read or write latency.
@@
Lastly, we expect LV [I] ’s global version log (technique #4) to have a noticeable but smaller impact on latency in the single-table YCSB setup. As seen in Fig. 5(a), it replaces the base table’s local version assignment, adding a 4ms overhead for reads and writes compared to the baseline. Further, even though the validation phase takes longer for writes, it more significantly affects the latency of the reads. In LV [I] ’s WriteSerializable mode, version assignment is skipped for reads, thereby shortening their latency. In summary, the global version log introduces overhead for all operations in a sequential, single-table setting. However, the WriteSerializable mode of LV [I] can mitigate this impact in such scenarios.
@@
Figure 6. YCSB when using Spark (Delta Lake) and LV [R].
[…]
[YCSB when using Spark (Delta Lake) and LV-R.]YCSB when using Spark (Delta Lake) and LV [R]. Spark shows small throughput improvements and more aborts for more concurrent clients. LV [R] committing after each operation shows a slightly higher throughput than Spark with no aborts and increases with more concurrent clients. When bundling all operations in a single transaction, LV [DL] shows a high throughput with no aborts but a decrease in throughput for more concurrent clients due to the required redos.
5.3. Concurrent Workloads

### v2

@@ 5.2.3 Isolation
Base OTFs: Serializability (Single-Query, Single-Table). OTF formats guarantee serializability limited to single-query and single-table transactions journals/pvldb/ArmbrustDPXZ0YM20 . Since OTFs only feature single-query and single-table transactions, all transaction operations are contained within a single log file, providing serializability or snapshot isolation isolation_databricks ; conf/cidr/0001KPDSZ23 . As a result, the execution order of all transactions equals a serial execution. Their operations can be ordered transitively to form a serial history between all operations on the same table.
Lemma 6
OTF guarantee a serial ordering of transactions and their operations:
$T_{1},T_{2}\in T,tr(T_{1})=tr(T_{2}),|T_{1}|=|T_{2}|=1:T_{1}<T_{2}\Rightarrow\forall op_{1}\in T_{1},op_{2}\in T_{2}:op_{1}<op_{2}$
LV[R]: Serializability for single tables. LV[R] extends Lemma 6 by adding atomic markers placed during a transaction’s start to put all transaction operations into a single file. Hence, all distinct transactions operating on the same table initially get a unique version and are ordered by the base OTF structure (Lemma 6).
Proof Sketch: Hypothesis. LV[R] guarantees serializability for single-table transactions.
Proof Sketch: Proof by contradiction. Assume LV[R] produces a non-serializable history for all single-table, multi-query transactions $T$. Then, at least two transactions $T_{1},T_{2}\in T$ can not be ordered. Since both transactions operate exclusively on the same table (single-table transactions), they must add their operations to the same metadata later. Because $T_{1}\not\sim T_{2}$ and OTFs always define an order of operations, we know from Lemma 6 that there exists a circular dependency:
$\exists a,b\in T_{1},\exists x,y\in T_{2}:a<x\cap b>y$ (5)
However, the structure of OTFs (compare section 2) implies that at least one transaction must have written multiple files containing distinct operations.
Proof Sketch: Conclusion. This is a direct violation of LV[R] commit protocol (overwrite a marker with a log file containing all operations). Hence, different log files must belong to different transactions, and Equation 5 does not apply. Hence:
Lemma 7
LV[R] guarantees serializability for multi-query, single-table transactions using its marker technique:
$T_{1},T_{2}\in T,tr(T_{1})=tr(T_{2}),|tr(T_{1})|=1:T_{1}<T_{2}\Rightarrow\forall op_{1}\in T_{1},op_{2}\in T_{2}:op_{1}<op_{2}$
@@
Table 3: Shortened Hermitage Results (multi-query) for Spark using Delta Lake (Spark[DL]) or Iceberg (Spark[IB]) and LakeVilla (LV). Anomalies described with ”-MT” are multi-table versions of the original scenario. We tested all anomalies and show the ones with the biggest differences here
@@
G0-MT (Write Cycles) - * - * , - -
G1a (Aborted Reads) , , - - -
G1b (Interm. Reads) , , - - -
G1c-MT (Circ. information flow) , * , * , - -
OTV (Obs. Txn. Vanishes) , , - - -
P4 (Lost Update) , , - - -
Concurrent Reads (Linearizability) - - , , -
prevents anomaly: - ; not prevented: , .
@@
LV[CT]: Repeatable Reads. LV[CT] removes Lemman6’s restriction on single tables ($|tr(T_{1})|=1$). Because it reuses the marker module (#1), we only prove that LV[CT] ensures that it creates a serializable order of transactions across tables. All other proofs are analogous to LV[R]’s for operations on the same table.
Proof Sketch: Hypothesis. LV[CT] provides repeatable reads across tables for multi-table, multi-query transactions.
Proof Sketch: Direct Proof. LV[CT] sets markers in each accessed table of a transaction. Hence, it reserves a set of table versions within the respective metadata layers. Hence:
$\forall T_{i}\in T,\forall tbl\in tr(T_{i}),\exists w_{(tbl,T_{i})}\in\mathbb{N}$ (6)
By applying Lemma 7 per table in $T$, we see that this orders all operations for each table. However, unlike LV[R], if a transaction accesses a new table, LV[CT] sets a new marker in the respective metadata layer. Given two concurrent multi-table transactions $T_{1},T_{2}\in T$, $T_{2}$ decides to access a new table tbltbl. In this scenario, we have two possibilities:
Case 1: $T_{1}$ did not access tbltbl yet. In this case, $T_{1}$ executed no operations on tbltbl beforehand. Hence, all of $T_{1}$’s operations can be ordered before or after $T_{2}$’s marker version on tbltbl, depending on the existing relation $T_{1}\sim T_{2}$ on another table. Hence, $T_{1}$ is isolated by default. There is no need for LV[CT] to readjust this situation.
Case 2: $T_{1}$ already accessed tbltbl. Then, a marker for $T_{1}$ in tbltbl already exists. When $T_{2}$ sets a new marker, this automatically forms the relation:
$T_{1}<T_{2}\Leftrightarrow w_{(tbl,T_{1})}<w_{(tbl,T_{2})}$ (7)
If there are previously executed operations in $T_{1}$ and $T_{2}$, we have three scenarios that may apply here:
Case 2.1: $(tr(T_{1})\cap tr(T_{2})=\{tbl\}$. Hence, there is no other table that $T_{1}$ and $T_{2}$ have both accessed yet. Their global relation is updated to $T_{1}<T_{2}$, and both can continue. On commit, all of $T_{1}$’s operations are logically ordered before $T_{2}$. Hence, $T_{2}$ must compare its changes to the respective logs and either commit or abort.
Case 2.2: $\forall z\in(tr(T_{1})\cap tr(T_{2})/tbl,w_{(z,T_{1})}<w_{(z,T_{2})}$. In this scenario, both transactions encountered each other in another table beforehand (If they are empty, case 2.1 applies). In these tables, $T_{1}$ and $T_{2}$ already have the newly created relation within tbltbl globally. Thus, LV[CT] automatically upholds this relation, and both transactions can continue normally.
Case 2.3: $\forall z\in(tr(T_{1})\cap tr(T_{2})/tbl,w_{(z,T_{1})}>w_{(z,T_{2})}$ Hence, LV[CT] observes two contradicting relations in different tables and triggers a marker shift. As LV[CT] can not rearrange the order within tbltbl, the marker shift frees all markers in all other tables and places a new marker. This effectively inverts the relation in such tables as described in Lemma 1. Hence, with the relation found in tbltbl and after a maker shift, LV[CT] guarantees:
$\forall z\in tr(T_{1})\cap tr(T_{2}),w_{(z,T_{1})}<w_{(z,T_{2})}$ (8)
This fixes the global relation between those two transactions. Since Cases 2.1, 2.2, and 2.3 ensured consistent ordering between transactions, LV[CT] provides an order among the tables.
However, LV[CT] is not serializable due to the marker shift: A transaction can never read dirty data due to the marker replacement during the commit. To prevent Non-repeatable reads, LV[CT] ’s marker shift never changes the read version of each table $r$. Despite a compatibility check during commit, phantom reads can occur: If a transaction requests a range of tuples, it usually accesses multiple data files in the OTF table. As LV[CT] only remembers previous reads, it can not predict any changes incorporated by a marker shift (if a transaction that was previously dependent and now became preceding) belongs to a range query. Hence, for LV[CT], the same query might look like a different operation that will work with the new changes incorporated.
Proof Sketch: Conclusion. Because LV[CT] defines a structure between dependent transactions in all scenarios but might incur phantom reads, it provides repeatable reads globally.∎
Lemma 8
LV[CT] guarantees repeatable reads globally using its conflict detection technique
LV[I]: Serializability. LV[I] provides the isolation level of serializability. Its global version log defines a total order between all committed transactions. Each transaction must enter its atomic validation phase and check its changes using R1 – R3.
Proof Sketch: Hypothesis. LV[I] guarantees serializability
Proof Sketch: Proof by contradiction. Assume LV[I] is not serializable. Then, two transactions $T_{1},T_{2}\in T$ can not be ordered sequentially ($T_{1}\not\sim T_{2}$). Since LV[I] takes version assignment to its global version log, both transactions received their read and write versions. Since $T_{1}$ is able to read $T_{2}$’s data, we know:
$\exists tbl\in tr(t_{1})\cap tr(t_{2}):w_{(tbl,T_{2})}<r_{(tbl,T_{1})}$ (9)
Thus, the global version log must have previously known $w_{(tbl,T_{2})}$. Hence, $T_{2}$ must have passed the validation phase, followed by committing in all its accessed tables before making its changes globally available. However, $T_{2}$ observed $T_{1}$’s changes, implying:
$\exists tbl_{2}\in tr(T_{1})\cap tr(T_{2}):r_{(tbl_{2},T_{2})}>w_{(tbl_{2},T_{1})}$ (10)
Hence, $T_{2}$ received a version $r_{(tbl_{2},T_{2})}$ larger than $w_{(tbl_{2},T_{1})}$ from the global version log. Hence, $T_{1}$ must have passed validation and committed before $T_{2}$. However, equations 9 and 10 are impossible to fulfill at the same time with the atomic validation phase of LV[I]. Transactions pass through it sequentially and end it by committing or aborting. Hence, there can not be a circular dependency between $T_{1}$ and $T_{2}$.
Proof Sketch: Conclusion. Due to the atomic design of the global validation phase, LV[I] guarantees that previously validated operations will be ordered before and after the current transaction. Similarly, as the global version log is a global layer for the entire LH, it defines relations between all transactions, giving Serializability.∎
Lemma 9
LV[I] guarantees serializability globally using its global version log:
$T_{1},T_{2}\in T:T_{1}<T_{2}\Rightarrow\forall op_{1}\in T_{1},op_{2}\in T_{2}:op_{1}<op_{2}$
Moreover, if we relax R1 using the ’slack’ parameter to a larger area to let read-only transactions pass validation without a check, we skip the check on the transactions’ read sets.
Lemma 10
LV[I] with a relaxed R1 guarantees snapshot isolation globally.
5.2.4 Durability of LakeVilla.
Similar to the original OTF protocols, the guarantees of object stores and a client successfully uploading all new files before generating the log file (e. g., compare journals/pvldb/ArmbrustDPXZ0YM20 ) achieve Durability for LV[R] and LV[CT]. Further, a transaction only updates LV[I]’s global version log if it successfully enters validation and all tables’ data files are written. Hence, LV[I] uploads all data files to the object store before committing and can rely on the object store guarantees to provide Durability.
5.2.5 Experimental Validation
@@
To provide further validation of LV’s guarantees and show the combined effect of its features, we experimentally check Spark and our prototype using LV[R], LV[R, CT], and LV[R, CT, I] for the anomalies described by Hermitage hermitage . Table 3 shows a selection of the results with additional multi-table versions of the same anomaly (marked with -MT). The original Delta Lake (Spark[DL]) and Iceberg (Spark[IB]) were included despite their restriction on single-query transactions, which led them to fail most multi-query anomalies. An exception is G0-MT (write cycles), avoided by their limitations and thus marked uniquely (G0-MT). In contrast, LV[R] prevents table-level anomalies like lost updates (P4), aborted reads (G1a), or Intermediate Reads (G1b) and adapts to aborted/crashed transactions (OTV). When adding LV[CT], the design can resolve conflicts across tables like Write Cycles (G0-MT) and Circular information flow (G1c-MT). Further, LV[I]’s global isolation guarantees concurrent transactions to observe globally consistent versions across tables. In summary, these results confirm our targeted guarantees.
6 Evaluation
@@
Overview. We start by analyzing the different phases of the LV Features and comparing them to Spark’s Delta Lake spark_web ; delta_lake_web . We investigate their effects under concurrent workloads using YCSB [^1] YCSB-LH available: https://github.com/goetztj/YCSB-LH conf/cloud/CooperSTRS10 ; YCSB-C and test different feature combinations based on the cloud analytics benchmark (CAB) journals/pvldb/RenenL23 . We conclude by running LV on selected analytical queries from TPC-DS, testing for long-term effects with LSTBench, and estimating the added overhead for systems like Spark.
6.1 Experimental Setup
@@
Object Store: AWS S3 and MinIO. For all experiments, we either use AWS S3 s3_web or MinIO minioWeb on a cloud instance with an Intel(R) Xeon(R) Platinum 8260 CPU with 48 cores at 2.40GHz with 524GB of RAM and 5TB of SSD storage running Linux. S3 and MinIO provide strict consistency minio_str_consistency ; aws_str_consistency and conditional writes aws_conditional_writes . Further, MinIO offers analytical features for the experiments. Both services are accessed through the AWS S3 SDK. An experiment uses S3 is not stated differently.
@@
Spark: Delta Lake. Spark 3.5.3 runs with Hadoop 2.12, Delta Lake 3.0.0 (checkpoint interval 10), and Hadoop AWS 3.3.4. We minimized caching to observe the OTFs’ effects on the system. Our HIVE 3.0.0 catalog and its PostgreSQL backend run in two separate docker containers. Spark uses Java 11 and runs distributed on 9 machines with the same setup as above. The coordinator runs on one server with specifications like our MinIO server and shares its resources with the catalog store and benchmark drivers to reduce the networking effects between these components. For workers, we configured two instances per server using an Intel(R) Xeon(R) Platinum 8260 CPU at 2.40GHz with 8 cores each. Each worker used 64GB RAM and eight cores.
6.2 Phases of OTFs
@@
Figure 7: Phases for read and write workloads of a single transaction for YCSB queries on MinIO (matplotlib)
Comparison of Spark and LakeVilla. We implemented our LV prototype as a separate C++ client. Thus, we first compare the differences between it and other systems. For this purpose, we ran YCSB sequentially with Spark using Delta Lake (Spark[DL]) and LV[DL]. If not stated otherwise, we use the default YCSB of 1000 operations on 1000 entries with row sizes of 1KB conf/cloud/CooperSTRS10 . We logged the different phases of both implementations and calculated both the average latency for reads and writes, and how much each phase contributes to the overall latency. We expect both systems to display similar behavior. We show the results in Figure 7. Spark[DL] and LV[DL] indeed behave similarly for reads. They spend most of their time accessing and reading the metadata of the YCSB table (95% and 93%, respectively) and the rest to load the requested data.
@@
The different LV Features. Next, we evaluate the impact of each LV feature on latency. We use the same setup as before and show our results in Figure 7(a). First, we analyze the overhead of placing LV[R]’s and LV[CT]’s markers (module #1). We hypothesize that it affects the latency, especially for reads. In Figure 7(a), writes show minimal impact as they align with metadata layer operations in the original OTF (LV[DL]). Further, freeing those markers is fused with the write transaction’s log placement, compensating for its effect as seen in Figure 7(a). In contrast, reads experience a 20% latency increase due to added OTF synchronization, with an additional impact when we free the markers post-read. This indicates that LV[R] and LV[CT] are better suited for write operations.
@@
Next, we analyze the impact of LV[R] ’s sublog (module #2) and LV[CT]’s deadlock detection (module #3) on latency using the same setup. Because both modules are expected to work concurrently alongside the standard operations, we expect minimal impact on latency, which is also confirmed by Figure 7(a). The time spent by the prototype on the actual operation in LV[R] and LV[CT] matches the baseline (LV[DL]). LV[R]’s sublog generation is distinct and does not impact the baseline operation. Further, LV[CT]’s deadlock detection does not detect problems as YCSB uses only one table. Hence, under non-concurrent workloads, neither module impacts read or write latency.
@@
Lastly, we expect LV[I]’s global version log (module #4) to have a noticeable but smaller impact on latency in the single-table YCSB setup. As seen in Figure 7(a), it replaces the base table’s local version assignment, adding a 4ms overhead for reads and writes compared to the baseline. Further, even though the validation phase takes longer for writes, it more significantly affects the latency of the reads. In LV[I]’s WriteSerializable mode, version assignment is skipped for reads, thereby shortening their latency. In summary, the global version log introduces overhead for all operations in a sequential, single-table setting. However, the WriteSerializable mode of LV[I] can mitigate this impact in such scenarios.
@@
Figure 8: YCSB when using Spark (Delta Lake) and LV[R] on S3 and MinIO (matplotlib)
6.3 Concurrent Workloads